OpenStudy (anonymous):

Ms. Rao buys a computer, a printer, and a scanner for $2,543. The computer cost $1,503 more then the printer. The printer cost $123 more than the scanner. How much does the computer cost? can someone please help me?

OpenStudy (nubeer):

scanner =x printer = x+123 computer = x+123+1503 2453 =scanner + pinter +computer .

OpenStudy (anonymous):

so the computer cost 2,453 or 1,626

OpenStudy (nubeer):

hmm actually none of it.. x+x+123+x+123+1503 = 2453

OpenStudy (nubeer):

find x that will be price of scanner and subsitute this x in above equations to find price of computer and printer.

OpenStudy (anonymous):

ok

OpenStudy (nubeer):

sorry *2543